### v2

# A comparison between best-fit eccentricity definitions and the standardized definition of eccentricity

Abstract
In the absence of a unique, gauge-independent definition of eccentricity in General Relativity, there have been efforts to standardize the definition for Gravitational-Wave astronomy. Recently, Shaikh et al. proposed a model-independent measurement of eccentricity $e_{\text{gw}}$ from the phase evolution of the dominant mode. Many works use loss functions (LFs) to assign eccentricity to a reference waveform, for instance by fitting a Post-Newtonian expression to assign eccentricity to Numerical Relativity (NR) simulations. Therefore, we ask whether minimizing common LFs on gauge-dependent model parameters, such as the mismatch $\mathcal{M}$ or the $L_{2}$-norm of the dominant mode $\mathpzc{h}_{22}$ residuals, for non-precessing binaries, ensures a sufficient $e_{\text{gw}}$ agreement. We use $10$ eccentric NR simulations and the eccentric waveform TEOBResumS-Dalí as the parametric model to fit on eccentricity $e_{0}$ and reference frequency $f_{0}$. We first show that a minimized mismatch, the $\mathcal{M}\sim 10^{-3}-10^{-2}$ results in better $e_{\text{gw}}$ fractional differences ($\sim 1\%$) than with the minimized $\mathpzc{h}_{22}$ residuals. Nonetheless, for small eccentricity NR simulations $(e_{\text{gw}}\lesssim 10^{-2}$), the mismatch can favor quasi-circular ($e_{0}=0$) best-fit models. Thus, with sufficiently long NR simulations, we can include $e_{\text{gw}}$ in the LF. We explain why solely fitting with $e_{\text{gw}}$ constitutes a degenerate problem. To circumvent these limitations, we propose to minimize a convex sum of $\mathcal{M}$ and the $e_{\text{gw}}$ difference to both assign non-zero eccentric values to NR strains and to control the mismatch threshold.
I Introduction
Gravitational Wave (GW) observations have opened up a new window to the universe. The ground-based network of GW observatories LIGO-Virgo Aasi et al. (2015); Acernese et al. (2015) has detected about 90 compact binary coalescences (CBCs) in the first three observing runs Abbott et al. (2023) that includes detection of binary black hole (BBH), black hole-neutron star (BHNS) Abbott et al. (2021a) and binary neutron star (BNS) Abbott et al. (2017, 2020) systems. One of the key questions that we hope to address using these detections is how these binaries form in nature, and the eccentricity of the binary orbit is one of the indicators that could play an important role in this regard.
Binaries lose orbital eccentricity due to the loss of energy and angular momentum via GW radiation Peters and Mathews (1963); Peters (1964). It is expected that for binaries formed via isolated evolution (e.g., in the galactic fields  Mapelli (2020)), the orbit of the binary will become circularized by the time the GW emitted from such CBCs enters the sensitive frequency band of the ground-based GW detectors. Indeed, so far all the detection and analyses of the CBC events by the LIGO-Virgo-KAGRA (LVK) Aasi et al. (2015); Acernese et al. (2015); Akutsu et al. (2021) collaboration has been performed assuming the binary to be in a quasicircular orbit Abbott et al. (2021b)). However, recent studies indicate that the GW from binaries formed in a dense environment via dynamical formation Samsing et al. (2022); Tagawa et al. (2021); Kozai (1962); Lidov (1962); Naoz (2016); Antonini et al. (2017); Randall and Xianyu (2018); Zevin et al. (2019); Yu et al. (2020); Bartos et al. (2023); Bhaumik et al. (2024) (e.g., in globular clusters, galactic nuclei, field triples etc.) could retain non-negligible eccentricity when entering the sensitive band of the ground-based GW detectors. Thus detection of non-negligible eccentricity could indicate the dynamical formation of the binary. For instance, debates as to the characteristics and eccentricity of GW190521, along with other merger events, are still ongoing Romero-Shaw et al. (2020); Gayathri et al. (2022); Romero-Shaw et al. (2021); Siegel et al. (2023); Gupte et al. (2024).
Consequently, the GW community has undertaken significant efforts to develop tools and frameworks for detecting and analyzing GW data from such binaries on eccentric orbits. Several groups have been actively developing gravitational waveform models for binaries in eccentric orbit using semi-analytic formalism like Effective-One-Body (EOB) and Post-Newtonian approximation as well as employing fully Numerical Relativity (NR) simulations Warburton et al. (2012); Osburn et al. (2016); Cao and Han (2017); Liu et al. (2020); Ramos-Buades et al. (2022a); Nagar et al. (2021); Islam et al. (2021); Liu et al. (2022); Memmesheimer et al. (2004); Huerta et al. (2014); Tanay et al. (2016); Cho et al. (2022); Moore et al. (2018); Moore and Yunes (2019); Van De Meent and Warburton (2018); Chua et al. (2021); Hughes et al. (2021); Katz et al. (2021); Lynch et al. (2022); Klein (2021); Wang et al. (2025); Shi et al. (2025); Phukon et al. (2025); Sridhar et al. (2024); Gamboa et al. (2024a, b); Morras et al. (2025); Planas et al. (2025). , Several works demonstrated the need to include eccentricity by showing potential bias in estimated parameters and spurious violation of General Relativity due to neglecting eccentricity Favata et al. (2022); O’Shea and Kumar (2023); Saini et al. (2022, 2023); Favata (2014); Shaikh et al. (2024); Das et al. (2024); Kumar et al. (2025). The inclusion of other physical effects of eccentric binaries, such as the argument of periapsis (e.g. Clarke et al., 2022), is also being discussed.
Modelling waveforms from compact binaries in eccentric orbits has one challenging aspect. Unlike Newtonian orbits, the binary orbit in a strong gravity regime can not be described by a closed ellipse – the orbit goes through pericenter precession. In addition, due to GW radiation, the size of the orbit and the eccentricity reduce towards merger time Zevin et al. (2019). There is no unique definition of eccentricity in General Relativity and, therefore, to capture features of a binary in an eccentric orbit, different waveform models adopt internal definitions as per the convenience of the model framework. This leads to an unwanted ambiguity regarding the eccentricity parameter of waveform models Knee et al. (2022). Such ambiguity could be reflected in the posteriors of parameter inference from GW strain data that include eccentricity, and would impede our understanding of astrophysical populations among research groups choosing different eccentric waveform models.
To avoid such ambiguity, there has been efforts to standardize the definition of eccentricity based on a gauge-independent quantity like the GW modes at infinity Ramos-Buades et al. (2020); Islam et al. (2021); Bonino et al. (2023); Ramos-Buades et al. (2022b); Shaikh et al. (2023). Since such a definition uses only the information in the GW modes, it removes the model dependency, i.e. it is independent of the way the GW modes are generated. This definition makes use of the oscillatory nature of the frequency and amplitude evolution of the eccentric GWs–it determines the local maxima and minima from waveform quantities like the frequency or the amplitude. Consequently, the computation requires a minimum (typically $\gtrsim 5$) number of orbits in the waveform to reliably estimate eccentricity. Thus far, the definition has been examined primarily for waveforms in which the $(2,2)$ multipole is dominant. See Ref. Shaikh et al. (2023) for more details on strategies to implement a standardized definition of eccentricity, which we refer to as $e_{\text{gw}}$. Recently, Patterson et al. (2025) proposed an eccentricity identification, also restricted to the $(2,2)$-mode, from an harmonic decomposition of the waveform and Boschini et al. (2025) developed a gauge independent estimation method from catastrophe theory. The authors of Islam and Venumadhav (2025) propose a publicly available framework for smooth, monotonic eccentricity estimates invoking universal modulation features of non-precessing binaries Islam (2024a); Islam and Venumadhav (2024).
Although advances in NR formalism and large computational resources have enabled NR groups to simulate increasingly long compact binary coalescences (CBCs), including eccentric ones, many NR simulations lack a sufficient number of orbits to reliably apply the standardized definition of eccentricity. For such shorter NR simulations, other definitions of eccentricity may be used to assign an eccentricity value to these simulations. One commonly used definition of eccentricity involves the minimization of a Loss Function (LF) by adjusting the parameters of a chosen waveform model to best match the given NR waveform. The mismatch $\mathcal{M}$ Flanagan and Hughes (1998)111or equivalently its counterpart to maximize, the overlap or faithfulness e.g. in Hinder et al. (2018), is commonly used in such procedure since this scalar quantity defines the optimal matched filter in the GW detection problem. The mismatch can either play the role of the LF to minimize (e.g. Joshi et al., 2023) or can be used as a validation metric after minimization of another LF. For instance, the authors of Habib and Huerta (2019) computed their LFs from the evolution of the dimensionless quantity $M\omega$, with $M$ the BBH total mass and $\omega$ the mean orbital frequency, and then calculated the mismatch to quantify the accuracy of the characterization of the eccentric NR simulation at hand.
Moreover, the mismatch is also routinely used as the LF to validate eccentric waveform models against NR simulations (e.g. Hinder et al., 2018; Liu et al., 2020; Nagar et al., 2021; Bonino et al., 2023). Other applications of the mismatch include mapping the eccentricity definitions between distinct waveform models, as Knee et al. (2022) extensively studied for the models TEOBResumS-Dali Chiaramello and Nagar (2020); Nagar and Rettegno (2021) and SEOBNRE (Liu et al., 2020), or developing fast surrogate models Blackman et al. (2015, 2017a, 2017b); Islam et al. (2021). For instance, Yun et al. (2021) fitted a reduced-order (RO) surrogate for SEOBNRE, covering eccentricities up to $\sim 0.25$ to improve GW inference computational capabilities. We stress that these works used the mismatch as a validation tool for the fitted surrogates, not to fit the surrogate models beforehand.
Alternatively, one can choose the $L_{2}$-norm, the distance between vectors of strain data in time domain as the LF to compare a model (waveform, surrogate model to train) from a set of NR simulations Blackman et al. (2017a); Islam et al. (2021, 2022); Islam (2024b); Rink et al. (2024). We will thus also include this $L_{2}$-norm in our analysis.
However, a minimized mismatch, by definition, makes sure that the waveform modes in frequency domain are best correlated, and does not necessarily guarantee an optimal agreement of the eccentricity evolution independently computed from the NR and the model waveform as per the standardized definition. This limitation motivates us to ask whether these two definitions are equivalent in practice, i.e. does the eccentricity evolution $e_{\text{gw}}$ independently computed from an NR waveform match the measured $e_{\text{gw}}$ of a model waveform parametrized so that it minimizes a LF? We address this question by taking a set of NR waveforms and computing the difference in the eccentricity between these two approaches. We consider NR waveforms that are long enough so that they contain enough number of orbits for the standardized definition to be applicable. Studying the case of TEOBResumS-Dalí for which eccentricity is defined with two parameters – the initial eccentricity $e_{0}\in[0,1.]$ and reference frequency $f_{0}$ – with numerical search strategies similar Knee et al. (2022), we find that there does not exist a simple one-to-one correspondence between the best match waveforms and the eccentricity evolution. A lower mismatch is not equivalent to a better $e_{\text{gw}}$ agreement, and conversely we find that fitting for eccentricity agreement on $e_{\text{gw}}$ results in degenerate solutions and is thus not sufficient to characterize eccentric NR simulations. In particular, we will interpret the previous points regarding the limitations of a model allowing to define eccentricity through $e_{0}$ and $f_{0}$ only, without the mean anomaly $l_{0}$. We will also discuss the very recent findings of Bonino et al. (2024) who used the standardized $e_{\text{gw}}$ definition to map eccentricity between NR simulations and waveform models. The authors of this work, proposed to determine new pairs of initial conditions based on the time difference of the first apastra for TEOBResumS-Dalí or similar EOB waveforms.
We organize the study as follows. In Section II, we introduce some useful reminders about the mathematical formalism for waveforms and about eccentricity in GW astronomy. The next section deals with the characterization of NR eccentric simulations using the mismatch and other LFs with a given waveform approximant. We incorporate the measurement of eccentricity as developed in Shaikh et al. (2023) in this process, and emphasize the limitations of the mismatch to define eccentric waveforms and then propose an alternative, fitting procedure, assuming we can only modify $(e_{0},f_{0})$. Finally, we propose a new, simple LF to circumvent aforementioned limitations and we discuss the implications of our findings when it comes to assessing the agreement between eccentric approximants when relying solely on mismatch.
II Theoretical Background
II.1 Gravitational waveform and mismatch
The plus ($h_{+}$) and cross ($h_{\times}$) polarizations of GWs can be combined to define a complex strain $\mathpzc{h}\equiv h_{+}-ih_{\times}$. The complex strain $\mathpzc{h}$ can be decomposed into a sum of spin-weighted spherical harmonics modes $\mathpzc{h}_{\ell m}$ along any direction $(\iota,\varphi_{0})$ in the binary’s source frame as
$\mathpzc{h}(t,\iota,\varphi_{0})=\sum_{\ell=2}^{\infty}\sum_{m=-\ell}^{\ell}\mathpzc{h}_{\ell m}(t)\prescript{}{-2}{Y}_{\ell m}(\iota,\varphi_{0})$ (1)
where $\iota$ and $\varphi_{0}$ are the polar and azimuthal angles, respectively, on the sky in the source frame and ${}_{-2}{Y}_{\ell m}$ is the spin = -2 weighted spherical harmonics.
The mismatch $\mathcal{M}$ between two given GW strains $\mathpzc{h}_{1}$ and $\mathpzc{h}_{2}$ is the bounded scalar defined as (Cutler and Flanagan, 1994)
$\mathcal{M}=1-\arg\max_{t_{c},\phi_{c}}\frac{\langle\mathpzc{h}_{1},\mathpzc{h}_{2}\rangle}{\sqrt{\langle\mathpzc{h}_{1},\mathpzc{h}_{1}\rangle\langle\mathpzc{h}_{2},\mathpzc{h}_{2}\rangle}},$ (2)
where $t_{c}$ and $\phi_{c}$ are the coalescence time and phase, respectively, and the inner product $\langle\mathpzc{h}_{1},\mathpzc{h}_{2}\rangle$ is
$\langle\mathpzc{h}_{1},\mathpzc{h}_{2}\rangle=4\mathcal{R}\int_{f_{\text{low}}}^{f_{\text{high}}}\frac{\tilde{\mathpzc{h}_{1}}^{\star}(f)\tilde{\mathpzc{h}}_{2}(f)}{S_{n}(f)}\mathrm{d}f.$ (3)
Here, $f$ is the frequency, $\tilde{\mathpzc{h}}_{1,2}$ are the Fourier transforms of the corresponding time domain waveforms, $\star$ represents the complex conjugate and $\mathcal{R}$ represents the real part. Further, $f_{\text{low}}$ and $f_{\text{high}}$ denote the lower and higher frequency cutoff of the detector band and $S_{n}(f)$ denote the one-sided power spectral density (PSD) of the detector noise. We recognize in equation (3) the inner product of square-integrable functions, weighted by $S_{n}$, and in the second term of equation (2) the correlation coefficient in $[-1,1]$ in Fourier space.
II.2 Eccentricity definitions
In the Newtonian regime, the orbital eccentricity can be uniquely defined using the orbital separations at apocenter ($r^{\text{a}}$) and pericenter ($r^{\text{p}}$)
$e_{\text{Newt}}=\frac{r^{\text{a}}-r^{\text{p}}}{r^{\text{a}}+r^{\text{p}}}\,,$ (4)
In the GR regime, the binary orbit is no longer a closed ellipse. Due to the GW radiation, the orbit loses energy and angular momentum, causing the orbital size to shrink as the system inspirals. In addition, GR effect induces pericenter precession. In the quasi-Keplerian formalism, these effects are captured by extending the Keplerian equation of motion through the introduction of different eccentricity parameters — the radial ($e_{r}$), temporal ($e_{t}$) and the angular ($e_{\phi}$) eccentricities. These quantities are usually defined in terms of the conserved energy and the angular momentum which are gauge dependent. In the EOB formalism, the initial conditions for the dynamics are prescribed in terms of an eccentricity parameter that is defined using the quasi-Keplerian formalism, making it a gauge dependent definition. Similarly, in the self-force calculations of extreme mass ratio inspirals, we characterize eccentricity using the turning points of geodesics which constitutes again a gauge dependent choice since it depends on underlying coordinate systems used to describe the background spacetime metric. Finally, for the NR simulations, we fit the compact object trajectories to analytical PN or Newtonian expressions to define eccentricity. Since the trajectories are dependent on the gauge used in the simulation, the subsequent eccentricity characterization is also gauge dependent.
II.2.1 Standardized eccentricity definition
To avoid the gauge ambiguity and model dependence, a definition of eccentricity based solely on the gravitational waveform has been proposed  Ramos-Buades et al. (2022a); Islam et al. (2021); Ramos-Buades et al. (2020); Bonino et al. (2023). This eccentricity $e_{\omega_{22}}$ is computed using the interpolants through frequency of the (2, 2) mode at the apocenter ($\omega_{22}^{\text{a}}(t)$) and pericenter ($\omega_{22}^{\text{p}}(t)$)
$e_{\omega_{22}}(t)=\left(\frac{\sqrt{\omega_{22}^{\text{p}}(t)}-\sqrt{\omega_{22}^{\text{a}}(t)}}{\sqrt{\omega_{22}^{\text{p}}(t)}+\sqrt{\omega_{22}^{\text{a}}(t)}}\right)\,,$ (5)
where frequency $\omega_{22}$ is obtained from $\mathpzc{h}_{22}$ using the following decomposition into amplitude $A_{22}$ and phase $\Phi_{22}$
$\mathpzc{h}_{22}$ $=A_{22}e^{-i\Phi_{22}}$ (6)
$\omega_{22}$ $=\frac{\mathrm{d}\Phi_{22}}{\mathrm{d}t}.$
However, $e_{\omega_{22}}$ does not have the correct Newtonian order limit. Therefore, to ensure the correct Newtonian limit, Shaikh et al. (2023) proposed a new waveform based definition of eccentricity $e_{\text{gw}}$.
$e_{\text{gw}}$ $=\cos(\Psi_{22})-\sqrt{3}\sin(\Psi_{22})$ (7)
$\Psi_{22}$ $\equiv\frac{1}{3}\arctan\left(\frac{1-e_{\omega_{22}}^{2}}{2e_{\omega_{22}}}\right)\,.$ (8)
Because $e_{\text{gw}}$ uses interpolants through frequency at the pericenters and apocenters, this definition requires the waveform to contain a minimum number of cycles ($\sim 5$) to build the interpolants robustly. Therefore, for shorter waveforms, we might not be able to reliably compute $e_{\text{gw}}$ and have to resort to other methods. In many NR simulations, this could be the case where the simulations have only a few final orbits of the compact object binary.
II.2.2 Eccentricity of the best fitted waveform
One alternative way of assigning an eccentricity to an NR waveform is to find the initial data ($f^{\text{best-fit}},e^{\text{best-fit}}$), where $e^{\text{best-fit}}$ is the initial eccentricity at the initial frequency $f^{\text{best-fit}}$ of the best-fit waveform,
[…]
where, the Loss Function (LF) $\mathcal{E}$ is a measure of disagreement between the NR and the model generated waveform $\mathpzc{h}_{\text{model}}$ at $(f_{0},e_{0})$. For example, one usually use the mismatch $\mathcal{M}$ (defined in equation (2)) as a good representation of $\mathcal{E}$. We discuss possible choices for LFs in Section III.3.2.
This method, unlike equation (7), uses an waveform model and therefore is model dependent. However, it does not depend on the number of orbits contained in the waveform and, therefore, can be applied to any waveform.
In the following sections, we test the robustness of the definition $e^{\text{best-fit}}$ in equation (9) for assigning eccentricity to NR simulations when compared to the $e_{\text{gw}}$ computed from the NR waveform. We accomplish this by studying a set of longer simulations where both of these methods are applicable.
III Methodology and setup
III.1 NR simulations and model waveform
For this study, we use $10$ eccentric simulations, as listed in Table 2, from the Simulating eXtreme Spacetimes (SXS) catalog (Boyle et al., 2019; SXS Collaboration, ; SXS, ) and run by Islam et al. (2021). We sort them in this table by $e_{\text{gw}}$ at $18$Hz from the measurements of Fig. 1. We use the implementation in the Python library gw_eccentricity  [^2] https://github.com/vijayvarma392/gw_eccentricity from Shaikh et al. (2023) to compute $e_{\text{gw}}$ in this work. We explain the chosen method later in Section III.2.
All the NR simulations in Table. 2 have mass ratio $q\equiv m_{2}/m_{1}=1$ ($m_{2}\leq m_{1}$). We convert the strain from geometric units to physical units by fixing the binary mass to $M=m_{1}+m_{2}=60M_{\odot}$ and its luminosity distance to $D_{L}=400$ Mpc similar to the parameters of GW150914.
To assign $e^{\text{best-fit}}$ to these simulations, we use the publicly available eccentric waveform model TEOBResumS-Dalí  [^3] https://bitbucket.org/teobresums/teobresums/. We have used the git commit 0f19532a9a1012b556bd4c72da10a1317d0e8751. To compute $e_{\text{gw}}$ and the LF $\mathcal{E}$, we will make use of the dominant (2, 2) mode only and output physical units strain from TEOBResumS-Dalí, In addition, since TEOBResumS-Dalí allows to define the reference frequency as the waveform frequency at apastron, periastron or the mean of the two, we choose the latter definition as in Knee et al. (2022). The inclusion of higher modes has notably been investigated by e.g. Habib and Huerta (2019); Bonino et al. (2023), and is beyond the scope of our study. It should also be kept in mind that the definition of the eccentric initial conditions are sensitive to changes in input parameters $e_{0},f_{0}\sim\mathcal{O}(10^{-6})$ Nagar and Rettegno (2021).
Table 1: Basic information about our set of equal mass, eccentric NR simulations from SXS. We analyze the equal mass binaries with physical units corresponding to total mass $M=60M_{\odot}$ and luminosity distance $D_{L}=400$Mpc.
Simulation ID $\lfloor N_{\text{orbit}}\rfloor$ $e_{\text{gw}}(18\mathrm{Hz})$
SXS:BBH:2294 $27$ 0.0028
SXS:BBH:2267 $26$ 0.011
SXS:BBH:2270 $26$ 0.027
SXS:BBH:2275 $26$ 0.059
SXS:BBH:2280 $26$ 0.083
SXS:BBH:2285 $24$ 0.103
SXS:BBH:2290 $21$ 0.139
SXS:BBH:2311 $27$ 0.168
SXS:BBH:2300 $24$ 0.168
SXS:BBH:2305 $25$ 0.198
III.2 Measuring $e_{\text{gw}}$
For $e_{\text{gw}}$ measurements, we consider the signals up to $2$ orbits before merger. These measurements rely, as currently implemented, on interpolants through $\omega_{22}$ at apastron and periastron in equation (5). The authors of Ref. Shaikh et al. (2023) propose different methods to locate apatron and periastron times, and for the sake of simplicity we will present only our choice for this work. We could not measure $e_{\text{gw}}$ from the straightforward Amplitude method, which locates extrema directly in $A_{22}(t)$, for small eccentricity cases. Instead, chose the ResidualAmplitude method instead that is more robust for small eccentricities at the cost of using a quasi-circular counterpart [^4] We found the computational cost to be lighter than with the AmplitudeFits or FrequencyFits methods also proposed in gw_eccentricity. to locate the extrema of $A_{22}(t)-A_{22}^{\mathrm{circ}}(t)$. The quasi-circular counterpart of TEOBResumS-Dalí will simply be the same waveform computed with $e_{0}=0$, while we will use the default IMRPhenomT approximation to measure $e_{\text{gw}}$ of our SXS data. Figure 1 shows the $e_{\text{gw}}$ measurements in our SXS set and emphasizes our motivation to prefer the ResidualAmplitude, in accordance to the conclusions of the original work about small eccentricities. Indeed, with the Amplitude method, we cannot measure $e_{\text{gw}}$ in a frequency range as large as with ResidualAmplitude for SXS:BBH:2294, SXS:BBH:2267 and SXS:BBH:2270, and the agreement ratio with ResidualAmplitude is the lowest in these cases (right panel of Figure 1).
Figure 1: Left: Frequency evolution of eccentricity $e_{\text{gw}}$ in the simulations from Table. 2. Right: $e_{\text{gw}}$ ratio between the ResidualAmplitude and Amplitude method from gw_eccentricity Shaikh et al. (2023). The former is more robust to the small eccentric NR strains in our set, justifying our preference. We resort to physical units by choosing a total binary mass $M=60M_{\odot}$ and luminosity distance $D_{L}=400$Mpc.
III.3 Methodology
III.3.1 Comparison study strategy
To test the closeness of $e^{\text{best-fit}}$, assigned to an NR simulation, when compared to $e_{\text{gw}}^{\text{NR}}$ measured directly from the NR waveform, we use the following strategy:
  1. 1.
    We compute $e_{\text{gw}}^{\text{NR}}$ directly from the given NR waveform.
  2. 2.
    Using a waveform model and a LF, we find the best-fit waveform $\mathpzc{h}^{\text{best-fit}}$ by finding the initial frequency $f^{\text{best-fit}}$ and eccentricity $e^{\text{best-fit}}$ by minimizing the LF using equation (9). In this work, we use TEOBResumS-Dalí, as mentioned in Section III.1, for generating the model waveform $\mathpzc{h}_{\text{model}}$ to find $\mathpzc{h}^{\text{best-fit}}$. In Section III.3.2, we discuss different LFs that can be used to do such a study.
  3. 3.
    Next, using equation (7), we compute $e_{\text{gw}}^{\text{best-fit}}$ of the best matched waveform $\mathpzc{h}^{\text{best-fit}}$ obtained in the previous step.
  4. 4.
    Finally, we compute the disagreement between $e_{\text{gw}}^{\text{NR}}$ and $e_{\text{gw}}^{\text{best-fit}}$
    $\Delta e_{\text{gw}}=e_{\text{gw}}^{\text{NR}}-e_{\text{gw}}^{\text{best-fit}}$ (10)
$\Delta e_{\text{gw}}$ acts as a measure of the disagreement of the eccentricities between the two methods of assigning eccentricity to a given NR simulation. One can also think of it as a measure of disagreement of the eccentricities of an NR waveform and the best-fit model waveform for a given waveform model.
III.3.2 Loss Functions
One of the crucial part of finding $(f^{\text{best-fit}},e^{\text{best-fit}})$ is to use an appropriate LF. In this work, we consider the following LFs:
  1. 1.
    Mismatch: The mismatch $\mathcal{M}$, defined in equation (2), between the NR and the TEOBResumS-Dalí waveform for a given initial frequency and eccentricity, is one of most used LF to measure the disagreement between two given GW waveforms. The mismatch also depends on the noise sensitivity $S_{n}(f)$ of the detector. Since the detectors are sensitive only within a particular window of frequencies, the mismatch only carries information from the relevant part of the signals given a particular detector, and is thus geared to practical applications rather than theoretical comparison of models. A "White Noise" mismatch, i.e. a flat $S_{n}(f)$, on the other hand, is suitable for theoretical comparisons of waveform modes below the Nyquist frequency for a given sampling rate. We considered two representative cases, namely the "White Noise" mismatch $\mathcal{M}_{W}$, i.e. a flat $S_{n}(f)$, and the aLIGO-like mismatch $\mathcal{M}_{L}$. We only emphasize analyses dealing with the former and show some insight about the $LF\sim\mathcal{M}_{L}$ in Appendix A.2, since the two different mismatch weightings led us to similar conclusions.
  2. 2.
    $L_{2}$-norm of $\mathpzc{h}_{22}$ residuals The match intervening in equation (2) is a correlation coefficient in Fourier space but not a distance, i.e. derived from a norm itself defined from the inner product of the vector space. One can therefore use a more direct measure of the differences between the two waveforms by computing in time domain555for time domain waveforms, as it it the case in the scope of this paper the normalized $L_{2}$ error
[…]
    $\equiv\sqrt{\|\Delta\mathpzc{h}_{22}\|/\|\mathpzc{h}_{22}^{\text{NR}}\|^{2}}\equiv\|\delta\mathpzc{h}_{22}\|.$
    We will note, for simplicity, $\mathcal{E}_{\mathpzc{h}_{22}}(f_{0},e_{0})=\|\delta\mathpzc{h}_{22}\|$, also emphasizing that this scalar is a distance in time-domain functions and not a correlation measure of Fourier space vectors as the mismatch. This LF is at the core of works such as Blackman et al. (2017a); Islam et al. (2021, 2022); Islam (2024b); Rink et al. (2024), notably to fit surrogate parametric models to reference NR simulations.
  3. 3.
    $\Delta e_{\text{gw}}$: In the cases where the waveforms are long enough, we can use $\Delta e_{\text{gw}}$, defined in equation (10), in the LF to obtain ($f^{\text{best-fit}},e^{\text{best-fit}}$). This is particularly important in finding the best matched model waveform for constructing a hybrid eccentric waveform, where the full inspiral-merger-ringdown waveform is built by smoothly stitching a model waveform (for inspiral part) and an NR waveform (for post-inspiral part). However, one should be careful about choosing the best matched waveform based solely on the the basis of minimal $\Delta e_{\text{gw}}$ without ensuring that this waveform also gives a reasonably small $\mathcal{M}$. Ideally the $\mathcal{M}$ associated with the minimized $\Delta e_{\text{gw}}$ should be $<1\%$. Therefore in Sec. IV.1.3, we will also address the question of whether minimizing $\Delta e_{\text{gw}}$ also ensures $\mathcal{M}<1\%$, without additional calibration.
  4. 4.
    $C_{\lambda}$: Finally, following the discussion in above, we study whether a combination of $\Delta e_{\text{gw}}$ and $\mathcal{M}$ would serve as a better LF than using only one of them to ensure both a small $\Delta e_{\text{gw}}$ and a small mismatch, $\mathcal{M}<1\%$.
    $C_{\lambda}=\lambda|\Delta e_{\text{gw}}|+(1-\lambda)\mathcal{M},$ (12)
    where $\lambda$ is a scalar value in the range $[0,1]$ and the $\Delta e_{\text{gw}}$ term can be measured on a single frequency bin or computed from the average difference over multiple bins in the inspiral phase.
III.3.3 LF minimization strategy
Figure 2: Duration of TEOBResumS-Dalí varying on $(e_{0},f_{0})$ with the fiducial physical units chosen in this work. The yellow bounds above indicate the durations, with the same physical units, of the shortest and longest SXS simulations at our disposal. Physical quantities correspond to total binary mass $M=m_{1}+m_{2}=60M_{\odot}$ and luminosity distance $D_{L}=400$Mpc.
An eccentric system requires two eccentricity parameters—eccentricity and mean anomaly at a given reference frequency—to accurately describe it. However, in our version of TEOBResumS-Dalí, the mean anomaly parameter is fixed to a certain value, and it is not a free parameter. Therefore, to effectively explore the same parameter space, we vary the reference frequency. These two approaches are equivalent since we have only two additional independent parameters to vary for eccentric systems (e.g. Knee et al., 2022). For all SXS waveforms, we estimate the duration of the inspiral as the time between the global maxima of $|A_{22}(t)|$ and the time at the start of the waveforms, for the physical units determined by our fiducial component masses and luminosity distance. We show in Figure 2 the waveform duration of TEOBResumS-Dalí in $(e_{0},f_{0})$ for the same physical units. We will use these durations to initialize the range of TEOBResumS-Dalí parameter searches.
As preliminary, we illustrate the need to vary both $e_{0}$ and $f_{0}$. While Joshi et al. (2023) fixed $f_{0}$ to fit waveform templates to NR sets by varying $e_{0}$ (in the context of studying the impact of higher $\ell$ modes), we show in Figure 3 that in our study case fixing $f_{0}$ to some values would not result in neither satisfying $e_{\text{gw}}$ agreements nor sub-percent mismatches in general. Indeed, we see in this plot that
  • we must also vary $f_{0}$ in addition to $e_{0}$ to minimize the mismatch (or any other LF a-priori) with some reference eccentric NR template.
  • In addition to $e_{\text{gw}}$ from our SXS set that gives us an approximate idea of the parameter search range, we must also use NR data to find an appropriate search range for $f_{0}$.
  • The noise PSD $S_{n}$ will have an impact on the best-match waveform. Even though this question was part of our initial motivations, we only show results using two different mismatches in Table 2, since our overall prescriptions are the same.
Figure 3: Minimization of the mismatch from equation (2), for three cases of fixed reference frequency $f_{0}$, between TEOBResumS-Dalí waveforms and some reference eccentric NR template, when $S_{n}$ is White Noise (left panel) or representative of aLIGO (right panel). Physical quantities correspond to total mass $M=m_{1}+m_{2}=60M_{\odot}$ and luminosity distance $D_{L}=400$Mpc. This example illustrates the need to perform 2D searches over $(f_{0},e_{0})$, rather than fixing $f_{0}$ and searching over $e_{0}$ in our case.
The most crucial step in obtaining the ($f^{\text{best-fit}},e^{\text{best-fit}}$) pair is finding an robust strategy to minimize the LF over ($f_{0},e_{0}$). This is because the LFs based on waveform modes have multiple local valleys making it hard to find the global minimum. We illustrate the shapes of $\|\delta h_{22}\|$ and $\mathcal{M}_{W}$ in the $(e_{0},f_{0})$ plane, determined for each SXS simulation metadata, in Figure 4.
Figure 4: LFs on coarse grid from SXS:BBH:2285. We indicate the coordinates of LF minima on the grid with cyan circles. Top panel: Mismatch with flat $S_{n}(f)$ referred to as $\mathcal{M}_{W}$. Bottom panel: $L_{2}$-norm of the $\left(2,|2|\right)$ mode residuals referred to as $\|\delta\mathpzc{h}_{22}\|$ (equation 11. Physical quantities correspond to total binary mass $M=m_{1}+m_{2}=60M_{\odot}$ and luminosity distance $D_{L}=400$Mpc.
In order to initialize the parameter search scope from NR data, one can use, as in Islam et al. (2021), estimates of the mean anomaly and eccentricity from the NR waveforms. Since our TEOBResumS-Dalí implementation allows to vary only eccentricity and reference frequency, we will use measures of $e_{\text{gw}}$ and the waveform durations instead.
  1. (1)
    For all SXS waveforms, we find appropriate values of the initial TEOBResumS-Dalí parameters by requiring the duration of the inspiral to be similar. Moreover, we have already measured $e_{\text{gw}}(f)$ in the left panel of Figure 1, which gives us appreciations of eccentricity range in the early part of the inspiral.
  2. (2)
    Next, we perform a coarse grid search over $(f_{0},e_{0})$ in the range determined by the previous step (1), as shown in Figure 4 for instance.
  3. (3)
    Find a small rectangular box around the rough parameters obtained in step (2) and use a Nelder-Mead search given some LF to get the refined pair $(f^{\text{best-fit}},e^{\text{best-fit}})$.
The coarse grid computation to initialize a Nelder-Mead search is in the spirit of Knee et al. (2022), although this work used a much thinner spacing on the waveform parameters coarse grid search. For this reason, at the cost of more brute force computations, we double-checked our $(e^{\text{best-fit}},f^{\text{best-fit}})$ pairs from step (3) by initializing $\mathcal{O}(100)$ Nelder-Mead searches at each $\tilde{e_{0}^{i}}$ that minimizes the LF for each frequency bin $f_{0}^{i}$ of the coarse grid (e.g. the vertical axes of both panels of Figure 4). While these redundant searches did not modify the conclusions, we obtained slightly lower LF values for some cases, thus allowing for more confidence in our minima.
In the next section, we comment our LF minimizations on our $10$ SXS simulations and study the correspondence with the measured $e_{\text{gw}}$.
IV LF minimization and eccentricity
To answer the aspects of this study, we first show the resulting $e_{\text{gw}}$ evolutions from the best-fit TEOBResumS-Dalí waveforms in Section IV.1. Then, we discuss the results and propose prescriptions to use $e_{\text{gw}}$ when fitting eccentric waveforms in Section IV.2. All the plots stem from the best-fit results detailed in Table 2, and the reader can jump to Figure 11 as a companion to our conclusions and prescriptions.
IV.1 Fitting waveforms with Loss Functions
As a starting point, Figure 5 presents the fractional $e_{\text{gw}}$ differences
$\delta e_{\text{gw}}(f)\equiv\left|\frac{e_{\text{gw}}^{\text{TEOB}}(f)}{e_{\text{gw}}^{\text{NR}}(f)}-1\right|,$ (13)
in frequency domain, during the inspiral phase, measured after finding the best-fit parameters $(e^{\text{best-fit}},f^{\text{best-fit}})$ for each LF introduced in Section III.3.2. We stress that the $e_{\text{gw}}$-based LF is calculated with the absolute difference $\Delta e_{\text{gw}}$ more appropriate as a LF especially for small eccentricity cases - while we will show the fractional difference of equation (13) in some plots. To ease the discussions, we will designate SXS:BBH:2294, SXS:BBH:2267 and SXS:BBH:2270 from Table 2 as the lowest eccentricity simulations from now on.
Figure 5: Fractional $e_{\text{gw}}$ differences from TEOBResumS-Dalí generated from $(e^{\text{best-fit}},f^{\text{best-fit}})$ with respect to each SXS simulation from Table 2. Each of the four panel corresponds to one LF of Section III.3.2. In the top left panel, the curves for SXS:BBH:2267 and SXS:BBH:2294 illustrate that the $\mathcal{M}_{W}$ minimization favors the quasi-circular regime of TEOBResumS-Dalí, resulting in $\delta e_{\text{gw}}\sim 100\%$ (up to numerical errors). The bottom right panel corresponds to $\lambda=0.3$. Physical quantities correspond to total mass $M=m_{1}+m_{2}=60M_{\odot}$ and luminosity distance $D_{L}=400$Mpc. Note that we only show a fraction of the measurable $e_{\text{gw}}$ frequency range, until $f=28$Hz, since the error $\delta e_{\text{gw}}(f)$ becomes less relevant as we get closer to merger time.
IV.1.1 LF = $\mathcal{M}$
In the top left panel of Figure 5, we observe that the highest $e_{\text{gw}}$ fractional differences correspond to to the low eccentricity cases. In particular, SXS:BBH:2294 and SXS:BBH:2267 are best-fit with $e_{0}=0$ using the mismatch. Except for SXS:BBH:2270, all best-match waveforms fall below $\delta e_{\text{gw}}\leq 5\%$ in fractional disagreement. Nonetheless, we will have to address two points in the next section.
  1. 1.
    the quasi-circular waveform can be enforced by the mismatch minimization for small eccentricities (in this case TEOBResumS-Dalí)
  2. 2.
    The mismatch and $e_{\text{gw}}$ disagreement are not one-to-one correspondent, i.e. the lowest mismatches do not result in better $e_{\text{gw}}$ agreements.
We note that using a different $S_{n}(f)$, in the aLIGO case, also led to a quasi-circular best mismatch, although not for the same cases (SXS:BBH:2294 in Table 2).
IV.1.2 LF = $\mathcal{E}_{\mathpzc{h}_{22}}$
The top-right panel of Figure 5 presents the $e_{\text{gw}}$ measurements with respect to SXS simulations after minimizing for equation (11), i.e. the $L_{2}$-norm in time domain of aligned $\left(2,|2|\right)$ modes. Overall, our minimization strategy from Section III.3.3 results in poorer $e_{\text{gw}}$ agreements than with the mismatch, especially for small eccentricities. We also note that this LF also resulted in a quasi-circular best-fit waveform in one case (SXS:BBH:2270) The poor agreements in comparison with the mismatch LF could be counter-intuitive since the $\left(2,|2|\right)$ mode residuals is a proper distance, whereas the mismatch is a correlation coefficient, i.e. a measure of the cosine of the angle between the vectors of Fourier coefficients "normalized" on the hypersphere. Moreover, as we will illustrate in more details in Section IV.2, $(e^{\text{best-fit}},f^{\text{best-fit}})$ from $\|\delta\mathpzc{h}_{22}\|$ minimization lead to relatively high mismatch best-fit waveforms, typically with $\mathcal{M}_{W}$ of a few percents (Table 2).
IV.1.3 LF = $\Delta e_{\text{gw}}$
In Figure 6, which shows $|\Delta e_{\text{gw}}|$ on the coarse $(e_{0},f_{0})$ grid, we observe that the valley extending in the $f_{0}$ direction is significantly narrower than in the case of the mismatch, although oscillations are still present. This valley also crosses local minima and maxima of the LFs of Figure 4. In the bottom-left panel of Figure 5, we show that as expected, the $e_{\text{gw}}$ agreement is the lowest when fitting directly on $\Delta e_{\text{gw}}$. What does not appear on this plot, however, is that the best-fit waveforms have high mismatches, up to $\sim 10\%$ (Table 2) for the highest eccentricity cases of Table 2. Thus, as one could expect from the fact the $e_{\text{gw}}$ only invokes the phase evolution $\Phi_{22}$ in equation (6), this LF is not appropriate on its own.
Figure 6: Illustration on SXS:BBH:2290 of a LF derived from $\Delta e_{\text{gw}}$, here $|\Delta e_{\text{gw}}(f_{m})|,f_{m}=18$Hz. Physical quantities correspond to total binary mass $M=m_{1}+m_{2}=60M_{\odot}$ and luminosity distance $D_{L}=400$Mpc.
IV.1.4 LF = $C_{\lambda}$
Our motivation for $C_{\lambda}$ is twofold. First, for long NR waveforms that have a sufficient number of orbits, a simple linear combination of two LFs, one of them computed with $e_{\text{gw}}$, is an easy solution to incorporate the standardized eccentricity definition. In addition, this hybrid LF allows by construction to deal with cases where mismatch minimization would enforce quasi-circular best-fit waveforms by controlling the level of $e_{\text{gw}}$ disagreement. While such cases primarily serve as a diagnosis of the limitations of the waveform in use, it is also important to characterize eccentricity in reference simulations with quantified $\mathcal{M}$ and $\Delta e_{\text{gw}}$ requirements which would, in turn, guide choices for $\lambda$. The case for finding the hyperparameter $\lambda$ is beyond the scope of our this paper, thus we will simply present results from particular values. Unless stated otherwise, we present results for $\lambda=0.3$, as in Figure, which allowed not to assign quasi-circular waveforms by minimization in the smallest eccentricity cases and to obtain best-fit waveforms with sub-percent mismatches. The bottom-right panel of Figure 5 demonstrates that the lowest eccentricity SXS cases are now fitted with an eccentric waveform with both sub-percent-mismatches (see Table 2 and Figure 11) and - except for the smallest eccentricity SXS:BBH:2294 - percent level best-fit $e_{\text{gw}}$ agreements.
IV.2 Interpretation and prescriptions
Figure 5, only tells us about the $e_{\text{gw}}$ agreement without actually informing about the similarity of the best-fit waveforms with their NR counterpart. Therefore, we show the mismatches of the best-fit waveforms for different LFs in addition to the $e_{\text{gw}}$ agreement evolution for two representative SXS examples in Figure 7, namely SXS:BBH:2280 and SXS:BBH:2311. As previously stated, fitting on $\Delta e_{\text{gw}}$ (orange lines) can generate best-fit waveforms with high mismatch values. Moreover, we indicate our $\mathpzc{h}_{22}$ minimization results and emphasize, in addition to the poorer $e_{\text{gw}}$ agreements, the higher mismatches $\mathcal{M}_{W}\sim 3\%$ in theses examples for this LF. On the left panel, for SXS:BBH:2280, the convex loss $C_{\lambda}$, at the cost of an additional mismatch of $\approx 2\times{10^{-4}}$, improves the $e_{\text{gw}}$ agreement from $\sim 3\%$ to $\sim 0.01\%$. Of course, the practical relevance of $C_{\lambda}$ would depend on thresholds for the mismatch given a waveform. On the right panel, minimizing $\mathcal{M}_{W}$ already gives a sub-percent $e_{\text{gw}}$ agreement in the measured band.
Figure 7: Frequency evolution of the fractional difference $\delta egw$ of the best-fit TEOBResumS-Dalí waveforms for our LFs from Section III.3.2, with respect to the measurements of the fixed SXS waveforms (one example in each panel.) The loss function $C_{\lambda}$ has $\lambda=0.3$ on the left panel and $\lambda=0.5$ on the right panel, where for instance $\lambda=0.3$ results in $|\delta e_{\text{gw}}|$ slightly above $1\%$ in this frequency range. Physical quantities correspond to total mass $M=m_{1}+m_{2}=60M_{\odot}$ and luminosity distance $D_{L}=400$Mpc.
Regarding our conclusion on LFs derived from $e_{\text{gw}}$, Figure 8 illustrates what we refer to as the "degeneracy" with the mismatch, which measures waveform similarity. As we previously emphasized, $e_{\text{gw}}$ considers only the phase evolution of the waveform. Additionally, its computation requires interpolation through the times of apastron and periastron. Therefore, an infinity of $(e_{0},f_{0})$ pairs can result in $|\Delta e_{\text{gw}}|$ below a required threshold. A difference of the order of $10^{-5}$ in $|\Delta e_{\text{gw}}|$ can either give a sub-percent mismatch or a mismatch of $\sim 30\%$ as in the example of Figure 8 where the carefully chosen extremes have opposite phases but the same interpolants through the local extrema of $\omega_{22}$.
Figure 8: Illustration on SXS:BBH:2300 of what we refer to as $e_{\text{gw}}$overfitting" when $LF\sim\Delta e_{\text{gw}}$. From an arbitrarily close $\sim\mathcal{O}(10^{-5})$ $e_{\text{gw}}$ fractional error, we explored the $(e_{0},f_{0})$ parameter space to find the worst and best mismatch possible, which correspond to $\omega_{22}$ interpolants having same opposite and same phase as the NR phase evolution, respectively. Physical quantities correspond to total mass $M=m_{1}+m_{2}=60M_{\odot}$ and luminosity distance $D_{L}=400$Mpc.
Figure 9: Time evolution of the real part of the $\left(2,|2|\right)$ mode of the best-fit TEOBResumS-Dalí waveforms for SXS:BBH:2290. The strains are coaligned, i.e. all the waveforms are shifted so that $t_{m}=0$ and $\Phi_{m}=0$ at merger time. Coalignment does not correspond to the time and phase shifts maximizing the match from equation (2). We indicate at the top of each panel, for each LF from Section III.3.2, the mismatch and the $e_{\text{gw}}$ difference of the best-fit waveform with the NR counterpart.
To explain our findings, particularly about the poor performances of the $L_{2}$-norm from equation (11), we present the time evolution of the $\left(2,|2|\right)$ mode for a specific SXS example, comparing it with the the best-fit TEOBResumS-Dalí waveforms in Figures 9 and 10. We chose to show in Figure 9 the coaligned waveforms rather than those shifted in time and phase as to maximize the match. Since all the strains have the merger time at $t=0$ and a null phase, this allows us to visually examine the actual computation for LF = $\mathcal{E}_{\mathpzc{h}_{22}}$ in time domain. While the red (SXS) and blue (TEOBResumS-Dalí) strains look the most similar in the upper-right panel, the mismatch is important ($\mathcal{M}_{w}=5\%$) and so is the $e_{\text{gw}}$ difference, as we can see in the earlier part of the inspiral on the plot. Figure 10 actually emphasizes the phase and amplitude evolution differences for the same $\left(2,|2|\right)$ mode as in the previous Figure 9. We argue that minimizing the difference of sampled $\mathpzc{h}_{22}$ vectors in time domain does not allow precise control over which part of the waveform is prioritized until merger nor does it ensure which frequency band should be trusted for a specific eccentric waveform model. There is no guarantee that minimizing equation (11) over a given time interval, on average, will ensure sufficient correlation of Fourier modes, as our examples suggest. This problem can be mitigated using the mismatch, though at the cost of relying on a specific $S_{n}(f)$.
Figure 10: Phase and amplitude evolution of the $\left(2,|2|\right)$ mode of the best-fit TEOBResumS-Dalí waveforms in the case of SXS:BBH:2290, where the waveforms have phase $\Phi_{m}=0$ at merger time $t_{m}=0$. This plot is a companion to Figure 9. Physical quantities correspond to total binary mass $M=m_{1}+m_{2}=60M_{\odot}$ and luminosity distance $D_{L}=400$Mpc.
Finally, In Figure 11, we sort the SXS simulations from Table 2 along the $x$-axis in order of increasing $e_{\text{gw}}$, and, when fitting TEOBResumS-Dalí to these simulations using either the mismatch or the mismatch combined $\Delta e_{\text{gw}}$ ($LF\sim\mathcal{M}_{W}$ or $LF\sim C_{\lambda}$, respectively), we display the corresponding fractional eccentricity error $\delta e_{\text{gw}}$ and mismatch from the TEOBResumS-Dalí waveform using $(e^{\text{best-fit}},f^{\text{best-fit}})$. On this Figure, we observe that all the $\mathpzc{h}^{\text{best-fit}}$ strains resulting from either of these two LFs have a mismatch $\mathcal{M}_{W}\leq 1\%$. Our fiducial case is $\lambda=0.3$; the case $\lambda=0.5$ led to mismatched higher than $1\%$ for the lowest eccentric cases and $\lambda=0.1$ resulted in a quasi-circular TEOBResumS-Dalí best-fit for SXS:BBH:2267. We argue that, in practical cases (e.g. fitting a surrogate model to a reference waveform or characterizing eccentric NR simulations with a model), the choice of $\lambda$ will be guided by specific match and eccentricity agreement requirements. Then, we emphasize that when fitting our TEOBResumS-Dalí implementation with only the mismatch for the small eccentricity SXS simulations (SXS:BBH:2294 and SXS:BBH:2267), the minimization procedure described in Section III.3.3 unequivocally favors the quasi-circular approximation $e_{0}=0$. The detailed values of Figure 11 can be found in the corresponding colored entries in Table 2.
Figure 11: Cross examination of the mismatch LF $\mathcal{M}_{W}$ with flat $S_{n}$ (equation (2), squares) and $C_{\lambda}$ LF with $\lambda=0.3$ ((12), circles). minimization results. Whichever LF is minimized, we show the mismatch of the best-fit TEOBResumS-Dalí waveform in red (left $y$-axis) and fractional $e_{\text{gw}}$ difference in indigo blue (right $y$-axis). The 10 SXS identifications on the $x$-axis are sorted by increasing $e_{\text{gw}}(f_{m}=18Hz)$ from Table 2. We emphasize that by using $LF\sim C_{\lambda}$, at the price of a higher mismatch than with $LF\sim\mathcal{M}_{W}$, depending on $\lambda$, we enforce eccentricity agreement by construction. For the smallest eccentric simulations SXS:BBH:2294 and SXS:BBH:2267, $C_{\lambda}$ allows to assign eccentricity to the NR strain rather than preferring the quasi-circular approximant with the mismatch.
Overall, we stress that we do not propose that $C_{\lambda}$ as a replacement for the mismatch, which ensures correlation of the waveform Fourier modes. Instead, $C_{\lambda}$ provides a way to address "failure" cases, such as when the model at hand favors a quasi-circular approximant, as observed for SXS:BBH:2294 and SXS:BBH:2267. Indeed, in the limit $\lambda\rightarrow 1$, we obtain an $e_{\text{gw}}$-based LF that can be overfitted, allowing the mismatch to increase arbitrarily between the worst- and best-matched waveform for a given order of magnitude in $\Delta e_{\text{gw}}$, as we illustrated in Figure 8. In other words, we advocate to use a very simple LF like $C_{\lambda}$ to deal with model limitations while controlling the mismatch threshold with $\lambda$, as emphasized in Figure 11.
V Conclusion and remarks
The recent development of the gauge-independent measurement $e_{\text{gw}}$ by the authors of Shaikh et al. (2023) motivated our study about LF minimization through numerical searches over the eccentricity parameters of a waveform model, with respect to a reference waveform. Our goal was to provide insight into the relationship between commonly used LFs, particularly the mismatch, and the resulting $\Delta e_{\text{gw}}$, i.e. the difference between the best-fit model and the fixed reference. Applications range from mapping eccentricity between waveform models, fitting fast surrogate models with the perspective or improving inference capabilities in the future, or characterizing eccentric NR simulations where the eccentricity is ill-defined in the GR regime from the initial conditions.
We conducted numerical searches to minimize LFs, with the strategy from Section III.3.3, on the eccentricity $e_{0}$ and reference frequency $f_{0}$ parameters of the EOB model TEOBResumS-Dalí with respect to $10$ eccentric SXS simulations with $e_{\text{gw}}$ ranging from $\sim 10^{-3}$ to $\sim 0.2$ at $18$Hz. Our fiducial LFs are the mismatch from equation (2) - "White Noise", i.e. with a flat $S_{n}(f)$ weighting -, the residuals in time domain of the $\left(2,|2|\right)$ mode $h_{22}$ as per equation (11), and LFs derived from the difference $\Delta e_{\text{gw}}$. As expected - since $e_{\text{gw}}$ utilizes only partial information from the dominant mode to determine interpolants through apastron and periastron times - fitting to arbitrarily small $e_{\text{gw}}$ differences can lead to high mismatches. Consequently, the best-fit waveform may not be physically meaningful as it does not represent the same binary system, even if the eccentricity evolution matches within the measured frequency range. Moreover, our strategy from Section III.3.3 with $h_{22}(t)$ residuals as the LF produces insufficiently matched waveforms ($\mathcal{M}\sim$ a few percents) and poor $e_{\text{gw}}$ agreements. We still report results for this LF in some plots and in the summary Table 2. When the LF is the mismatch, we can find $(e^{\text{best-fit}},f^{\text{best-fit}})$ pairs such that $\mathcal{M}\sim 10^{-3}$ for most of the SXS set and closer to $\sim 10^{-2}$ for the highest eccentricity cases ($e_{\text{gw}}\sim 0.2$). Moreover, the fractional difference $\delta e_{\text{gw}}$ remains within a few percents for $e_{\text{gw}}<0.2$. Nonetheless, for our two less eccentric SXS simulations, minimizing $\mathcal{M}$ can give quasi-circular $e_{0}=0$ best-fit waveforms. To circumvent this limitation - which may arise not only with TEOBResumS-Dalí, but also more generally - we propose a linear convex combination of the mismatch and a (positive) LF derived from $e_{\text{gw}}$. The parameter $\lambda$ controls the allowed deviation from the best $\mathcal{M}$, which, in the ideal case, represents the LF we would recommend based on our findings for the aforementioned practical applications.
In addition, the recent study Bonino et al. (2024) would corroborate this prescription on the mismatch, although not in the context of numerical searches as in our work. The authors, after assigning eccentricities to their own NR simulations set through a mapping procedure of the initial conditions, find $\mathcal{M}\sim 10^{-3}-10^{-2}$ (Figure 12 of current version) as a robustness metric rather than a LF. We argue that $C_{\lambda}$ is merely an alternate LF to deal with both parametric model limitations and mismatch requirements. A simple to implement LF like $C_{\lambda}$ will be particularly useful in the context of fitting surrogates to reference strains, among the practical applications that motivated our study. We leave new developments on fitting procedures invoking $e_{\text{gw}}$ in addition to classic LFs, with numerical searches or other strategies as in Bonino et al. (2024), to future studies.
Acknowledgements.
We thank Vijay Varma and Leo C. Stein for useful discussions. This work was supported by the National Research Foundation of Korea under grant No. NRF-2021M3F7A1082056.
Appendix A Additional Details
A.1 LF computations
Below, we provide some additional details about our computations of the LFs introduced in Section III.3.2 for the curious reader.
  1. 1.
    We estimate the mismatch expressed in equation (2) with the psd.filter.matchedfilter.match method from the publicly available library PyCBC Usman et al. (2016). The lower integration bound is $\sim f_{0}+0.02$Hz for the waveform to start slightly before the band and the higher integration bound is the Nyquist frequency determined the the sampling rate in physical units. The advanced LIGO approximation in the $\mathcal{M}_{L}$ cases - only shown in Table 2 - is the aLIGOZeroDetHighPower model from pycbc.psd
  2. 2.
    To compute the $L_{2}$ loss in time domain from equation (11), we coalign the waveforms and compute the numerical counterpart to equation (11) from the shared earliest time sample of the waveforms until merger time, which is defined by the maximum of the amplitude modulus $|A_{22}(t)|$. To do so, we shift the waveform and NR data in time and phase so that they have both the same phase $\Phi_{m}=0$ at merger time $t_{m}=0$. Then, to compute the LF using the same time vector for $\mathpzc{h}_{22}^{\text{NR}}(t)$ and $\mathpzc{h}_{22}^{\text{TEOB}}(t)$, we rely on linear interpolants.
  3. 3.
    To calculate $\Delta e_{\text{gw}}(f)$, we use linear interpolants in the common range $[f_{min},f_{max}]$ within which we can measure $e_{\text{gw}}$ for the waveform-NR pair. We compute either the difference measured at the specific bin $f_{m}=18$Hz, where we can measure $e_{\text{gw}}$ consistently in all our cases, or the average difference $\langle|\Delta e_{\text{gw}}|\rangle$ on frequency bins in the range $[f_{min},20Hz]$ where $f_{min}$ is the lowest frequency at which $e_{\text{gw}}$ is definite after $f=f^{\text{best-fit}}$.
A.2 LF minimization results
We present in Table 2 below a summary of the different LFs introduced in Section III.3.2 minimized with the strategy of Section III.3.3, including $C_{\lambda}$ from equation (12). The colored rows correspond to Figures 11 for reference, where $\lambda=0.3$.
Table 2: Characterization of SXS simulations with TEOBResumS-Dalí using different LFs, along with standardized $e_{\text{gw}}$ measurements. Blue and red entries correspond to Figure 11. The leftmost column indicates the chosen LF, and the "Values" are computed from the TEOBResumS-Dalí waveform that minimizes this LF. From left to right, the SXS:BBH:XXXX IDs are sorted by increasing $e_{\text{gw}}$ from Table 2. Reference frequencies $f_{0}$ in Hz and $f_{m}=18$Hz correspond to total mass $M=m_{1}+m_{2}=60M_{\odot}$ and luminosity distance $D_{L}=400$Mpc. Cases where the best-fit model has $\delta e_{\text{gw}}\sim 100\%$ correspond to a preferred quasi-circular model by that LF, which occurs when $LF\sim\mathcal{M}$ for the small eccentricity NR simulations.
LF Values
$\mathcal{M}_{W}$ $\mathcal{M}_{W}$
$e_{0}$
$f_{0}$
$\delta e_{\text{gw}}(f_{m})$
$\mathcal{M}_{L}$ $\mathcal{M}_{L}$
$e_{0}$
$f_{0}$
$\delta e_{\text{gw}}(f_{m})$
$\|\delta h_{22}\|$ $e_{0}$
$f_{0}$
$\delta e_{\text{gw}}(f_{m})$
$\mathcal{M}_{W}$
$|\Delta e_{\text{gw}}(f_{m})|$ $\delta e_{\text{gw}}(f_{m})$
$e_{0}$
$f_{0}$
$\mathcal{M}_{W}$
$\langle|\Delta e_{\text{gw}}|\rangle$ $e_{0}$
$f_{0}$
$\delta e_{\text{gw}}(f_{m})$
$\mathcal{M}_{W}$
$C_{\lambda}$ $e_{0}$
$f_{0}$
$\delta e_{\text{gw}}(f_{m})$
$\mathcal{M}_{W}$
2294 2267 2270 2275 2280 2285 2290 2311 2300 2305
0.22% 0.24% 0.45% 0.43% 0.67% 0.48% 0.34% 0.54% 0.74% 0.77%
0.000 0.000 0.027 0.073 0.103 0.130 0.162 0.226 0.218 0.262
14.203 13.958 14.070 14.041 14.280 14.921 16.098 14.514 14.936 15.201
99.89% 99.97% 21.23% 4.56% 2.83% 1.52% 0.22% 1.52% 0.59% 1.16%
0.62% 0.66% 0.92% 0.81% 0.96% 0.86% 0.81% 0.92% 1.02% 1.12%
0.000 0.029 0.041 0.072 0.096 0.133 0.159 0.225 0.216 0.260
14.203 14.032 14.070 13.690 14.259 14.128 16.083 14.511 14.925 15.188
99.89% 108.32% 21.51% 7.99% 9.48% 1.75% 1.59% 0.98% 0.50% 0.31%
0.012 0.011 0.000 0.035 0.083 0.109 0.145 0.228 0.218 0.253
13.690 13.352 12.400 13.044 13.494 14.481 15.537 13.802 14.146 15.136
240.90% 20.54% 100.00% 57.13% 25.45% 16.51% 13.01% 2.78% 4.48% 2.16%
1.61% 1.25% 0.94% 2.36% 2.77% 3.83% 5.02% 3.67% 4.04% 2.56%
3.0E-07 1.5E-07 1.1E-07 1.3E-07 1.1E-07 2.3E-07 8.4E-07 3.2E-07 7.2E-07 2.3E-06
0.003 0.015 0.034 0.080 0.116 0.118 0.181 0.222 0.214 0.249
13.670 13.265 14.107 13.343 13.135 16.087 14.535 14.528 15.105 15.711
1.59% 1.32% 0.50% 1.20% 7.76% 10.91% 20.54% 3.80% 26.39% 23.70%
0.002 0.015 0.031 0.084 0.102 0.119 0.154 0.258 0.203 0.265
14.666 13.101 15.231 12.836 14.755 15.994 16.836 12.699 15.839 14.901
0.12% 0.09% 0.08% 0.01% 0.00% 0.02% 0.00% 0.01% 0.00% 0.01%
1.45% 1.04% 3.30% 3.26% 3.22% 7.42% 14.13% 13.01% 4.47% 25.15%
0.001 0.014 0.034 0.076 0.106 0.128 0.162 0.225 0.217 0.261
14.271 14.024 14.079 14.047 14.288 14.912 16.094 14.509 14.930 15.195
8.81% 0.61% 0.20% 0.04% 0.01% 0.06% 0.06% 1.05% 0.29% 0.94%
0.22% 0.28% 0.46% 0.45% 0.69% 0.52% 0.34% 0.60% 0.76% 0.82%
For illustration, we also show the White Noise PSD mismatch $\mathcal{M}_{W}$ and aLIGO PSD $\mathcal{M}_{L}$ mismatch for SXS:BBH:2270, on the $(e_{0},f_{0})$ grid, in Figure 12. In the White Noise case (left panel), TEOBResumS-Dalí would actually prefer a quasi-circular computation $e_{0}=0$ to minimize the mismatch. The minimum would be different with the aLIGO model, which stresses the importance of choosing a representative PSD if one solely minimizes the mismatch from equation 2 to characterize an eccentric simulation. Note that a more representative PSD to target observational applications rather than theoretical does not guarantee to solve model limitations, as the $\mathcal{M}_{L}$ LF also favored a quasi-circular TEOBResumS-Dalí waveform in one case (SXS:BBH:2294, in Table 2).
Figure 12: Mismatch based LFs with SXS:BBH:2267 as the NR simulation, for two different PSDs $S_{n}$ (equation 3). Left panel: Mismatch $\mathcal{M}_{W}$ with White Noise PSD. Right panel: Mismatch $\mathcal{M}_{L}$ With aLIGOZeroDetHighPower analytical PSD model from Usman et al. (2016). Physical quantities correspond to total mass $M=m_{1}+m_{2}=60M_{\odot}$ and luminosity distance $D_{L}=400$Mpc.